OpenStudy (anonymous):

Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? f(n) = 28(0.5)n f(n) = 56(0.5)n-1 f(n) = 56(0.5)n f(n) = 112(0.5)n-1 please help?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

y=mx+b ->slope intercept form of the line. m=slope, b=y-intercept , (x,y) any point. first, find the slope. choose any two points (n, f(n)) from the given and use the slope formula. Do you know the slope formula?

OpenStudy (anonymous):

Y1-Y2 over X1-X2 right?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

I used the points (1, 56) and (2,28)

OpenStudy (anonymous):

@DLS

OpenStudy (anonymous):

@dumbcow can you please help me??

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

Is it B @greenlegodude57

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@driftracer305

OpenStudy (dumbcow):

just plug in n=1 and see which one gives you 56

OpenStudy (anonymous):

so its C @dumbcow

OpenStudy (dumbcow):

nope that gives 28

OpenStudy (anonymous):

oh its D i got C and D confused so am i right?

OpenStudy (dumbcow):

well no D gives 112 ... seems we are playing a guessing game. you now have a 50/50 shot

OpenStudy (anonymous):

I thought it was D because i divided 112 by .5 and got 56 so is it B then?

OpenStudy (dumbcow):

the exponent is "n-1" remember that initial value is n=1 1-1 = 0 ...... .5^0 = 1 thats why its 112 for D yes B is correct because .5^(1-1) = .5^0 = 1

OpenStudy (anonymous):

ok thank you do you think you could help me with another? @dumbcow